OpenStudy (anonymous):

If g(x) = 3(x − 2), find the value of x if g(x) = 6. 12 2.6 8 4

OpenStudy (anonymous):

g(x) = 3(x − 2), and g(x) = 6. g(x)=3x-6 6=3x-6 6+6=3x 12=3x 12/3=x 4=x